• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Shabestari

    Log entry time 08:08:36 on April23,2008

    Entry number 233903

    keyword=Owl Shift Summary

    Owl shift summary April 23, 2008

    SL : Mitra Shabestari
    TO : Mitra Shabestari
    3rd : Khem Chirapatpimol


    Summary
    ========

    pi0 Production : Kinematics I Part(III)

    BigBite = 43.5 degrees, BigBite current = 388.1 A
    LHRS = 20.5 degrees, LHRS momentum = 0.9600 GeV/c
    RHRS = 105.0 degrees, RHRS momentum = 0.4596 GeV/c


    ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

    00:00 Shift started. We are going to Controlled access.
    There was a problem with ROC9 and ROC10.
    Yi, swing shift's SL, has requested controlled access to hard reboot the ROCs.

    00:15 ----> We are in controlled access. Khem is going down to the hall to reboot the ROC.

    00:24 ----> Khem is back from hall and we are back to beam permit.
    Yi Qiang is working on his shift summary.

    00:32 ----> MCC called we asked for 5 uA on LH2 target. Continuing to take production data.

    00:34 ----> Run# 3653

    00:40 ----> We had high rate and ~70% deadtime. Bryan called.
    Threshold on the back and front wire chambers were zero. Set both of them to 5V. We are back to production for now.

    00:47 ----> Run# 3655 ; events = 908 K.....First run after hall access....HV tripped.

    01:06 ----> HV on all E, delta E, and Drift Chambers tripped due to beam trip.

    01:17 ----> Run# 3656 ; events = 1 M.....ROC5 has "TDCs active error condition" message. We are going to hareboot ROC5 at the end of this run.

    01:42 ----> Run# 3657 ; events = 1 M.....Good run after ROC5 hareboot.

    02:02 ----> Run# 3658 ; events = 1 M

    02:21 ----> Run# 3659 ; events = 1 M

    02:40 ----> Run# 3660 ; events = 1 M

    02:51 ----> Run# 20739 ; started Right HRS

    02:55 ----> BPMA Y was ~0.48 mm, asked MCC to adjust it to nominal value of 0.7 mm.

    02:57 ----> All BPMs look fine.

    03:00 ----> Run# 3661 ; events = 1 M

    03:19 ----> Run# 3662 ; events = 1 M

    03:38 ----> Run# 3663 ; events = 1 M

    03:58 ----> Run# 3664 ; events = 1 M.....ROC5 has "TDCs active error condition" message. We are going to hareboot ROC5 at the end of this run.

    04:18 ----> Run# 3665 ; events = 1 M

    04:39 ----> Run# 3666 ; events = 1 M

    04:56 ----> Run# 20740 ; started Right HRS

    05:00 ----> Run# 3667 ; events = 1 M

    05:19 ----> Run# 3668 ; events = 1 M

    05:41 ----> Run# 3669 ; events = 1 M

    05:57 ----> Run# 3670 ; events = 1 M

    06:20 ----> MCC called. They have vacuum problem. No beam for a while.

    07:37 ----> MCC called. They are still working on vacuum problem, and will not be able to deliver beam for (at least) one hour.

    08:00 ----> Owl Shift ended with no beam.




    Beam enery before MCC's vacuum problem was 1192.39 MeV.
    For the current kinematics ( I ) at the end of the owl shift: Total Charge = 0.313 C